Given a recent outbreak of illness caused by E. coli bacteria, the mayor in a large city is concerned that some of his restaurant inspectors are not consistent with their evaluations of a restaurant's cleanliness. In order to investigate this possibility, the mayor has five restaurant inspectors grade (scale of O to 100) the cleanliness of three restaurants. The results are shown in the accompanying table. (You may find it useful to reference the q table.) Restaurant Inspector 72 68 73 69 75 84 85 80 82 84 54 59 60 56 xc a-1. Construct an ANOVA table. (Round intermediate calculations to at least 4 decimal places. Round "SS", "MS, "p-value" to 4 decimal places and "F" to 3 decimal places.) Answer is complete and correct. ANOVA Source of Variation Rows Columns Error Total df MS p-value 8.9330 1,723.6000 67.0670 1,799.6000 2.2330 0.8920 0.0000 0.266 28618000 102.799 8.3830 14
a-2. At the 5% significance level, can you conclude that the average grades differ by restaurant? Yes, since the p-value associated with restaurant is less than significance level. OYes, since the p-value associated with restaurant is greater than significance level. No, since the p-value associated with restaurant is less than significance level. No, since the p-value associated with restaurant is greater than significance level. b. If the average grades differ by restaurant, use Tukey's HSD method at the 5% significance level to determine which averages differ (If the exact value for n is not found in the table, use the average of corresponding upper& lower studentized range values. Negative values should be indicated by a minus sign. Round your answers to 2 decimal places.) Answer is not complete Population Mean Differences Does The Average Differ? Confidence Interval 8.18
Answer #1
column data , i (Factor B) 1 2 3
count, ni = 5 5 5
mean , x̅ i =Σxi / ni 71.400 56.800 83.000
Level of significance 0.05
no of treatments 3
df error 8
MSE 8.383
q-statistic value 4.043

critical value = q*√(MSE/n) = 5.2351

confidence interval = mean difference ± critical value

if confidence interval contans zero, then means are not different.

confidence interval
population mean difference critical value lower limit upper limit result
µ1-µ2 14.6 5.2351 9.36 19.84 yes,means are different
µ1-µ3 -11.6 5.2351 -16.84 -6.36 yes,means are different
µ2-µ3 -26.2 5.2351 -31.44 -20.96 yes,means are different